OpenStudy (anonymous):

Can you help me solve? 2/3x -1+ 1/6x=9

hero (hero):

\[\frac{2}{3x-1} + \frac{1}{6}x = 9\] Is that it?

hero (hero):

We haven't even started yet

OpenStudy (anonymous):

Sorry, go on.

hero (hero):

I was asking you if what I posted was the correct problem that you're working on.

OpenStudy (anonymous):

2 over 3 with an x on the side -1 + 1 over 6 with an x on the side =9. New to this web, not sure how to write out the problems. Can u understand what I am trying to say?

OpenStudy (anonymous):

2/3x -1 + 1/6x =9